OpenStudy (anonymous):

find dy/dx for y= cos (5/x)

OpenStudy (anonymous):

y=cos(5/x) dy/dx=-sin(5/x) .d/dx(5/x) do it further

OpenStudy (anonymous):

[5sin(5/x)] / x^2

OpenStudy (anonymous):

Thanks =]

geerky42 (geerky42):

Just use chain rule.